OpenStudy (anonymous):

[Help me :3 Reward: Medal and Fan] A bag of sweets contain 4 caramels, 3 fruit centres and 5 mints. If one sweet is taken out, what is the probablity that it is not a fruit centre? Options : a) 5/12 b) 3/4 c) 1/4 d) 1/3

OpenStudy (anonymous):

@rohitmathew Help meh bro?

OpenStudy (lxelle):

What do you think your choice of answr will be?

OpenStudy (anonymous):

5/12? @LXelle

OpenStudy (lxelle):

No, question says not fruit centres so 12-3=9

OpenStudy (lxelle):

Probability of getting 0 fruit centres.

OpenStudy (lxelle):

so 12/12 - 3/12 =?

OpenStudy (anonymous):

Oh yeah, thnx a lot @LXelle :)

OpenStudy (lxelle):

Pleasure's mine. :)

OpenStudy (anonymous):

ok..what i think is.. the total number of sweets is 12.. and it should not be a fruit centre. remainng sweets= 9 probability that it is something else=9/12=3.4

OpenStudy (anonymous):

3/4 sorry

OpenStudy (anonymous):

Thanks man @rohitmathew

OpenStudy (anonymous):

:)